OpenStudy (anonymous):

solve the equation 2sin^(2)x=2+cosx

OpenStudy (anonymous):

i was thinkng plugin 1-cos^2 but theres no theta symbol

OpenStudy (anonymous):

2sin^2(x) = 2 + cos(x) 2sin^2(x) = 1-cos(2x) cos(x)+cos(2 x)+1 = 0 cos(x)+cos(2x) = -1 x = pi/2, 2pi/3, 4pi/3, 3pi/2 for 0 <= x < 2pi

OpenStudy (anonymous):

2(1-cos^2x) = 2 + cosx cosx + cos^2x = 0 x = pi,pi/2

OpenStudy (anonymous):

2*sin^2x=2+cosx or, 2* (1-cos^x)=2+cosx (since sin^2x+cos^2x=1) or 2-2*cos^2x=2+cosx or 2*cos^2x+cosx=0 or cosx (2cosx+1)=0 thus cosx=0 or cosx=-1/2 so, x=pi/2 or 5*pi/3

OpenStudy (anonymous):

wait guys the x isnt part of the the exponent

OpenStudy (anonymous):

its sin^2 seperate then x